solve the following equation.

2log x = log 2x


1
Expert's answer
2021-09-06T14:42:00-0400

"2log x = log 2x"


"log x^2= log 2x" since "(xlogy= logy^x)"


Since the bases of logarithm are the same, set the arguments equal.


"x^2=2x"


"x^2-2x=0"


"x(x-2)=0"


"x=0"

"x-2=0"

"x=2>0."

Answer: "x=2."